The area of a rectangle is 384 square feet. If the perimeter is 80 feet, find the length and width of the rectangle.
Answer by swincher4391(1107)

A = lw
P = 2l + 2w
384 = lw
80 = 2l + 2w

384/l = w
80 = 2l + 2*(384/l)
80l/l = (2l^2+ 2*384)/l
80l = 2l^2 + 768
2l^2-80l+768 = 0
2(l^2-40l + 384) = 0
l = 16, l =24
Case 1: l=16
If l =16 then w = 384/16 = 24
If l =24 then w = 384/24 = 16
So, we could have 16x24 or 24x16.